# LED: Light Enhanced Depth Estimation at Night

### v2

@@ 5.5 Robustness Across Domains
We evaluate model robustness when operating beyond the original training domain. To this end, we run on HB data the model trained on the pattern and vice versa. We report results in Tab. 4.
@@
Furthermore, we demonstrate that incorporating only 10% of HB data is enough to mitigate the performance drop (see Fig. 6 (c)), allowing a unique network to perform well on both domains. In this case, the performance decline is reduced to +17.11% RMSE.
This adaptability allows a wider range of applications for the method. It enables selective use of the pattern when higher precision is necessary and its absence when avoiding glare is crucial. The pattern can be applied to specific regions such as objects of interest.
5.6 Real World Scenarios
We show the applicability of our method in real world scenarios using our in-house dataset. We report results in Tab. 5. LED significantly boosts performance over LB, across all metrics. In particular, DepthFormer (0.5°) shows -23.3% RMSE improvement. The qualitative result Fig. 7 shows LED’s robustness under interfering light sources, e.g., car headlights, street lights.
To account for objects passing through the pattern illumination, we investigate the impact of smaller checkerboard cells. Using a size of 0.25°, we observe great enhancement across all metrics (-17.4% RMSE, -18.8% Abs Rel against 0.5°). In contrast, further increasing the checkerboard resolution to 0.125° does not lead to significant improvements. Indeed, pattern features become difficult to distinguish with the camera resolution, thus reducing cues provided to the models. Performances shown by DepthFormer (LED) underpin the benefits of LED in complex real-world scenarios.
@@ 6 Limitations and Future Works
While the implicit learning of a reference pattern allows the method to be architecture agnostic, it remains limited to a single pattern defined during training. Moving forward specialized architectures handling the pattern as an integral input element would enable dynamic pattern optimization and improve method generalization.
We demonstrated LED capabilities on different projector-camera distance by training on both real (70 cm) and synthetic (150 cm) setup, meaning that it can be used over a same car model fleet. But it could be sensitive to large discrepancies. We intend to further study the generalization of LED in future works.
It is possible that LED pattern would glare other road users. One way to avoid this would be to leverage HD headlight to mask road users [39]. Safety issues and the impact of selective illumination on performance should be assessed in future works.
@@ 7 Conclusion
We introduce LED, a novel method for enhancing nighttime depth estimation, leveraging HD light patterns projected by modern vehicle headlights. Through extensive experiments, we demonstrate the effectiveness of our approach, showcasing notable improvements in depth perception within and beyond illuminated areas. The versatility of our method is further emphasized by its successful integration with two SOTA architectures: Adabins and DepthFormer. Moreover, LED demonstrated promising real-world performances. We release the Nighttime Synthetic Drive Dataset, a synthetic nighttime dataset comprising 49,995 comprehensively annotated images. This dataset will serve as a valuable resource for the research community, fostering exploration of various tasks related to autonomous driving in low-light conditions. As we tackle the difficulties of nighttime computer vision, our study highlight the potential advantages of using HD patterns for robust depth estimation.
Acknowledgments
We would like to thank Amandine Brunetto for her help and insights.
@@ Acknowledgments
A ROI Definition
B ROI-Only Training
To assess our method’s impact, we focus on the illuminated area. Thus, it is reasonable to evaluate performance when trained exclusively within the ROI. We report results of this experiment in Tab. S1. We observe that training solely within the ROI enhances Abs Rel and Sq Rel metrics, although other metrics show a decline. Since the network’s ability to estimate depth beyond the ROI is valuable for many applications, we opt not to pursue this approach.
Image input RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Table S1: Comparison of encoder-decoder performances between ROI-only training and baseline.
C Resolution Impact
Resolution and performance are usually highly correlated in computer vision, particularly in low-light conditions. To better understand the impact of resolution on our method, we train the encoder-decoder with center-cropped area resized at various resolutions. Results in Tab. S2 indicate that increasing resolution up to $640\times 640$ px does not improve performance. Conversely, decreasing resolution to $200\times 200$ px appears to enhance relative metrics while degrading others. Since most valuable cues in our method come from the pattern, these findings suggest that a resolution of $320\times 320$ px offers a favorable trade-off for pattern visibility and scene interpretation within our setup (as shown in Fig. S1).
Resolution RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
200 px 5.9413 0.1073 0.0206 0.1270 12.6227 3.9736 0.9545 0.9228 0.9921
320 px 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
640 px 5.7050 0.2109 0.0185 0.1420 14.1092 17.4085 0.9668 0.9837 0.9893
Table S2: Encoder-decoder performances at various resolutions.
Figure S1: Performances using multiple resolutions
D Encoder-Decoder Details
@@
LED uses a single pattern implicitly learned by the model, making it architecture-agnostic. This characteristic is demonstrated in the main paper (see Sec. 5.4) by applying LED to multiple state-of-the-art architectures [4, 23]. To prove our concept and conduct experiments, we opt for an encoder-decoder architecture with skip connections [34]. The detailed architecture is shown in Fig. S2. Despite the simplicity of this architecture, we show that LED enables the network to achieve performances comparable to or better than other tested SOTA architectures trained without our method.
Figure S2: Detailed encoder-decoder architecture
E Loss Study
@@
To determine the impact of each individual loss, we conduct an ablation study (see Tab. S3). Both L1 and Log L1 losses are tested for $\mathcal{L}_{depth}$ and $\mathcal{L}_{grad}$. Our findings indicate that each selected loss positively influences performance. The combination outlined in the paper (Sec. 3.3) demonstrates the most favorable trade-off between metrics.
@@
L1 Log L1 L1 Log L1 RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
@@
Table S3: Performance comparison using several combination of losses: 1st best, 2nd best, 3rd best.
F Examples of Other Patterns
To find a suitable pattern for our application, we generate half-sized datasets featuring commonly used structured light patterns: checkerboard, horizontal, and vertical lines (see Fig. S3). Comprehensive metrics from this experiment are available in Tab. S4, the checkerboard pattern demonstrates significantly superior performances.
Figure S3: Example of tested illuminations: high-beam (top-left), checkerboard (top-right), horizontal lines (bottom-left) and vertical lines (bottom-right)
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
@@ G LED Impact On Other Tasks
LED lighting pattern is mostly suitable for geometric tasks. We study its impact on other tasks performances, taking the example of semantic segmentation. To this end, we trained Mask2Former [10] using the available annotation in the Nighttime Synthetic Drive Dataset. Results are reported in Tab. S5. LED doesn’t improve, nor degrade performances. We even note a better stability with less variance over runs. Therefore, the similar performance on both domains suggests that LED enhances geometric tasks with limited impact on semantic ones.
Model mIoU $\uparrow$
Mask2Former (LB) 51.45 $\pm$ 4,00
Mask2Former (LED) 51.49 $\pm$ 0,76
Table S5: Semantic segmentation results on the Nighttime Synthetic Drive Dataset using Mask2Former [10]
H Real World Scenarios
To collect our real-world dataset, we made a prototype on a real car. We used an IDS U3-36L0XC camera and an Ouster OS1-128 LiDAR (Rev 7). Regarding the HD headlight, we opt for the Digital Micromirror Device technology, which offer the greatest resolution ($<$0.015°) with a horizontal FOV of 14° and vertical FOV of 7°. All the hardware was mounted on the roof of the car. This novel setup allowed us to test LED under another projector-camera configuration. Ground truth annotation was made using LiDAR data, aggregated and densified thanks to Exwayz software [12]. Due to the high resolution of the headlight we were able to collect data with smaller cell size than in simulation, giving more insight into pattern resolution impact. Model’s performances on the dataset demonstrate LED capabilities on complex real-world scenarios. We showcase more qualitative results in Fig. S4. A video of Depthformer (LED 0.25°) qualitative results on our real-world dataset is available in the supplementary material.
I Multiple Resolutions Robustness
Any HD headlight can project LED’s checkerboard, although with a resolution adjustment. To evaluate its impact on model performances, we propose to train a model with 3 checkerboard resolutions (0.5°, 0.25° and 0.125°) evenly split, and evaluate on each one. Table S6 demonstrates that the model is able to perform well on all resolutions although with a minor performance degradation compared to single resolution training. This shows our model’s ability to generalize over multiple resolutions, thus, various HD headlights or car models.
Test set RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$downarrow RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
@@ I Multiple Resolutions Robustness
Table S6: DepthFormer results when trained on 33% of each checkerboard resolutions (0.5°, 0.25°, 0.125°).
@@
We showcase the depth estimation enhancement obtained by comparing results with and without pattern. They are obtained by taking the encoder-decoder, trained on high beam (HB), and with our HD pattern (LED). Red boxes emphasize on enhanced regions. Some are zoomed in for improved visibility. LED results (right) exhibit higher precision, showcasing more accurate object boundaries and shapes compared to HB (middle). Far away obstacles are better defined and less blurry (first row), vehicles and pedestrians are sharper (second rows).
Figure S5: Zoomed in qualitative results showcasing LED improved depth prediction compared to high-beam (HB).
K Safety and Regulation
While using readily available hardware, LED is a research project not meant to be deployed on cars right away. Future works should assess potential safety issues. To maximize their safety, autonomous vehicles needs multiple perception mechanisms that ensure redundancy, compensating for blind spots of other sensors on the car, e.g., LiDAR in rain conditions. One can imagine that this feature could be turned off in crowded areas, or when detecting an incoming car after it was initially detected at a longer distance, ensuring safety. Regarding regulation, HD headlight technology is novel and the European regulation has authorized just recently the projection of specific HD pattern onto the road. Thus, laws are moving in this direction and evolving with the technology.

### v3

@@ 5.5 Robustness Across Domains
5.6 Real-world Scenarios
[…]
To account for objects passing through the pattern illumination, we investigate the impact of smaller checkerboard cells. Using a size of 0.25°, we observe great enhancement across all metrics (-17.4% RMSE, -18.8% Abs Rel against 0.5°). Further increasing the resolution to 0.125° does not lead to significant improvements.
[…]
In contrast, Depth Anything V2 achieves its best performance with a 0.125° resolution, as it is better suited to capturing fine details. When fine-tuned with only 1,000 images, we observe -24.3% RMSE compared to the LB baseline. Performances shown by DepthFormer and Depth Anything V2 underpin the benefits of LED in complex real-world scenarios.
@@ 6 Limitations and Future Works
The method’s reliance on a single reference pattern ensures agnosticism but limits its flexibility. Designing specialized architectures that treat the pattern as an input could enable dynamic pattern optimization and improve generalization.
We tested varying projector-camera distances (real: 70 cm, NSDD: 150 cm), showing LED robustness on similar car models. Future work will assess its usage with larger variations.
LED patterns may cause glare for other road users. Leveraging [Tamburo et al.(2014)Tamburo, Nurvitadhi, Chugh, Chen, Rowe, Kanade, and Narasimhan] to mask them could mitigate this. Future work should assess both safety concerns and the impact on performance.
@@ 7 Conclusion
We introduce LED, a method for enhancing nighttime depth estimation by leveraging high-definition light patterns projected by modern vehicle headlights. Through extensive experiments, we demonstrate significant improvements in depth perception, both within and beyond illuminated areas. Our method’s versatility is highlighted by its successful integration with two state-of-the-art architectures, Adabins and DepthFormer, as well as the foundation model Depth Anything V2. Moreover, LED shows promising real-world performance. We also release the Nighttime Synthetic Drive Dataset, comprising 49,995 fully annotated images. We hope it will serve as a valuable resource for the research community, supporting exploration of various nighttime perception tasks.
Acknowledgment
We would like to thank Amandine Brunetto for her help and insights. We also thank Naimeric Villafruela from the Fablab of Mines Paris for his assistance in creating the real-world prototype. In addition, we extend our warm thanks to the Exwayz team for the excellent collaboration on the real-world ground truth processing. This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011015334 made by GENCI.
@@ Acknowledgment
[…]
Mines Paris - PSL University
[…]
Paris, France
LED: Light Enhanced Depth Estimation at Night
A Safety and Regulation
While using readily available hardware, LED is a research project not meant to be deployed on cars right away. Future works should assess potential safety issues. To maximize their safety, autonomous vehicles needs multiple perception mechanisms that ensure redundancy, compensating for blind spots of other sensors on the car, e.g., LiDAR in rain conditions. One can imagine that this light pattern could be turned off in crowded areas, or when detecting an incoming car after it was initially detected at a longer distance, ensuring safety. Regarding regulation, HD headlight technology is novel and the European regulation has authorized just recently the projection of specific HD pattern onto the road. Thus, laws are moving in this direction and evolving with the technology.
B Encoder-Decoder Details
B.1 Encoder-Decoder Architecture
@@
LED uses a single pattern implicitly learned by the model, making it architecture-agnostic. This characteristic is demonstrated in the main paper (see section section˜5.1) by applying LED to multiple state-of-the-art architectures [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. To prove our concept and conduct experiments, we opt for an encoder-decoder architecture with skip connections [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. The detailed architecture is shown in fig.˜S1. Despite the simplicity of this architecture, we show that LED enables the network to achieve performances comparable to or better than other tested SOTA architectures trained without our method.
Figure S1: Detailed encoder-decoder architecture.
B.2 Learning Objectives
To train the encoder-decoder, we adopt a combination of loss functions inspired by [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani]. The primary loss, $\mathcal{L}_{depth}$, is the Log L1 loss presented in eq.˜1. It measures the error between the estimated depth, $d_{i}$, and the corresponding ground truth, $g_{i}$. This variant of L1 loss attributes less significance to errors occurring at greater distances. This adjustment aligns with the expectation that given errors, e.g., 1 m, should have greater impact when within a few meters of the camera but are more tolerable at extended distances:
$\mathcal{L}_{depth}=\frac{1}{N}\sum_{i=1}^{N}|log(d_{i})-log(g_{i})|.$ (1)
We address edges fidelity, particularly challenging in low-light conditions. To this end, we incorporate a loss, $\mathcal{L}_{grad}$, that specifically emphasizes on gradients. As we believe edges sharpness is important regardless of the distance, we employ a standard L1 Loss expressed in eq.˜2 instead of a logarithmic one used in [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani]. Moreover, our experiments revealed no performance improvements with the Log version. $\nabla_{x}(d_{i})$ and $\nabla_{y}(d_{i})$ respectively represent the spatial derivative of $d_{i}$ along the x and y-axis,
$\mathcal{L}_{grad}=\frac{1}{N}\sum_{i=1}^{N}|\nabla_{x}(d_{i})-\nabla_{x}(g_{i})|+|\nabla_{y}(d_{i})-\nabla_{y}(g_{i})|.$ (2)
Similar to [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani], we want to ensure accurate surfaces representation in depth maps.
The depth normals are estimated at each pixel using $n_{i}^{a}\equiv[-\nabla_{x}(a_{i}),-\nabla_{y}(a_{i}),1]^{\intercal}$. The cosine similarity loss, expressed in eq.˜3, is then employed to compare estimated and ground truth normals. $\langle.,.\rangle$ denotes vector inner product operation:
$\mathcal{L}_{normal}=\frac{1}{N}\sum_{i=1}^{N}|1-\frac{\langle n_{i}^{d},n_{i}^{g}\rangle}{\sqrt{n_{i}^{d},n_{i}^{d}}\sqrt{n_{i}^{g},n_{i}^{g}}}|.$ (3)
Finally, our learning objective can be expressed as eq.˜4, we set $\lambda_{1}=1$ and $\lambda_{2}=1$:
$\mathcal{L}=\mathcal{L}_{depth}+\lambda_{1}\mathcal{L}_{grad}+\lambda_{2}\mathcal{L}_{normal}.$ (4)
B.3 Loss Study
@@
To determine the impact of each individual loss, we conduct an ablation study (see table˜S1). Both L1 and Log L1 losses are tested for $\mathcal{L}_{depth}$ and $\mathcal{L}_{grad}$. Our findings indicate that each selected loss positively influences performance. The combination outlined in section˜B.2 demonstrates the most favorable trade-off between metrics.
@@
L1 Log L1 L1 Log L1 RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
@@
Table S1: Loss study. Performance comparison using several combination of losses: 1st best, 2nd best, 3rd best.
C Depth Anything V2 Details
To fine-tune Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] on our datasets, we used the metric model pre-trained on KITTI. We follow their indications for specializing the model to metric depth estimation. For each given number of training images, the split was selected randomly in the whole dataset. Figure˜S2 illustrates the RMSE of Depth Anything V2 across different numbers of fine-tuning images. We observe that the model is near its performance peak with just 1,000 images. All LED-enhanced models outperform the LB, also this architecture benefits from finer resolution in the pattern, demonstrating its best performances with LED 0.125°.
Figure S2: Few-shot learning performances of Depth Anything V2 on our real dataset. All LED patterns demonstrate superior performances over LB. The model is near its maximal performances with only 1,000 training images.
D ROI Details
D.1 ROI Definition
D.2 ROI-Only Training
To assess the impact of our method, we focus on the illuminated area. Thus, it is reasonable to evaluate performance when trained exclusively within the ROI. We report results of this experiment in table˜S2. We observe that training solely within the ROI enhances Abs Rel and Sq Rel metrics, although other metrics show a decline. Since the network’s ability to estimate depth beyond the ROI is valuable for many applications, we do not pursue this approach of ROI-only training.
Image input RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
Table S2: Encoder-decoder performances on NSDD between ROI-only training and reference from section section˜5.1.
Resolution RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
200 px 5.9413 0.1073 0.0206 0.1270 12.6227 3.9736 0.9545 0.9228 0.9921
320 px 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
640 px 5.7050 0.2109 0.0185 0.1420 14.1092 17.4085 0.9668 0.9837 0.9893
Table S3: Encoder-decoder performances on NSDD at various resolutions.
Figure S3: Encoder-decoder performances on NSDD using multiple resolutions
Figure S4: Example of illuminations pattern tested: high-beam (top-left), checkerboard (top-right), horizontal lines (bottom-left) and vertical lines (bottom-right).
E Resolution Impact
Resolution and performance are usually highly correlated in computer vision, particularly in low-light conditions. To better understand the impact of resolution on our method, we train the encoder-decoder with center-cropped area resized at various resolutions. Results in table˜S3 indicate that increasing resolution up to $640\times 640$ px does not improve performance. Conversely, decreasing resolution to $200\times 200$ px appears to enhance relative metrics while degrading others. Since most valuable cues in our method come from the pattern, these findings suggest that a resolution of $320\times 320$ px offers a favorable trade-off for pattern visibility and scene interpretation within our setup (as shown in fig.˜S3).
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
@@ E Resolution Impact
F Examples of Other Patterns
To find a suitable pattern for our application, we generate half-sized datasets featuring commonly used structured light patterns: checkerboard, horizontal, and vertical lines (see fig.˜S4). Comprehensive metrics from this experiment are available in table˜S4, the checkerboard pattern demonstrates significantly superior performances.
@@ G LED Impact On Other Tasks
We introduce the LED lighting pattern for improving performance on geometric tasks. We study its impact on the performance of other tasks that might be running in the same time on the vehicle, taking the example of semantic segmentation. To this end, we train Mask2Former [Cheng et al.(2021)Cheng, Misra, Schwing, Kirillov, and Girdhar] using the available annotation in the Nighttime Synthetic Drive Dataset. Results are reported in table˜S5. LED doesn’t improve, nor degrade performances. We even note a better stability with less variance over runs. Therefore, the similar performance on both domains suggests that LED enhances geometric tasks with limited impact on semantic ones.
Pattern mIoU $\uparrow$
[…]
LB 51.45 $\pm$ 4,00
LED 51.49 $\pm$ 0,76
Table S5: Mask2Former [Cheng et al.(2021)Cheng, Misra, Schwing, Kirillov, and Girdhar] Semantic segmentation results on NSDD.
Test set RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
@@
Table S6: DepthFormer results on real dataset when trained on 33% of each checkerboard resolutions (0.5°, 0.25°, 0.125°).
H Using Multiple Checkerboard Resolutions
Any HD headlight can project LED’s checkerboard, although with a resolution adjustment. To evaluate its impact on model performances, we propose to train a model with 3 checkerboard resolutions (0.5°, 0.25° and 0.125°) evenly split, and evaluate on each one. Table˜S6 demonstrates that the model is able to perform well on all resolutions although with a minor performance degradation compared to single resolution training. This shows our model’s ability to generalize over multiple resolutions, thus, various HD headlights or car models.
I Real World Scenarios
To collect our real-world dataset, we made a prototype on a real car. We used an IDS U3-36L0XC camera and an Ouster OS1-128 LiDAR (Rev 7). Regarding the HD headlight, we opt for the Digital Micromirror Device technology, which offer the greatest resolution (<<0.015°) with a horizontal FOV of 14° and vertical FOV of 7°. All the hardware was mounted on the roof of the car. This novel setup allowed us to test LED under another projector-camera configuration. Ground truth annotation was made using LiDAR data, aggregated and densified following DOC-Depth method [de Moreau et al.(2025)de Moreau, Mathias, Hassan, Yasser, Andrei, Hafid, Fabien, et al.] and thanks to Exwayz software [Exwayz(2024)]. Due to the high resolution of the headlight we were able to collect data with smaller cell size than in simulation, giving more insight into pattern resolution impact. The performance of our model on this dataset demonstrate LED capabilities on complex real-world scenarios. We illustrate additional qualitative results in fig.˜S5. A video of DepthFormer (LED 0.25°) qualitative results on our real-world dataset is available in the supplementary material.
Figure S6: Zoomed in qualitative results on NSDD showcasing LED improved depth prediction compared to high-beam (HB).
@@ I Real World Scenarios
We illustrate in fig.˜S6 the improvements of depth estimation achieved by our method by comparing results with and without pattern. They are obtained by taking the encoder-decoder, trained on high beam (HB), and with our HD pattern (LED). Red boxes emphasize on enhanced regions. Some are zoomed in for improved visibility. LED results (right) exhibit higher precision, leading to more accurate object boundaries and shapes compared to HB (middle). Far away obstacles are better defined and less blurry (first row), vehicles and pedestrians are sharper (second rows).